# A flexible quantum data bus

### v1

@@ Abstract
We consider multi-path routing of entanglement in quantum networks, where a pre-prepared multipartite entangled 2D cluster state serves as a resource to perform different tasks on demand. We show how to achieve parallel connections between multiple, freely chosen groups of parties by performing appropriate local measurements among diagonal paths, which preserves the entanglement structure of the remaining state. We demonstrate how to route multiple Bell-states along parallel lines via crossings, turns and fade-in/-outs, analogously to a data bus. The results apply to networks at any scale.
@@
Connecting devices constitutes the foundation of modern society. All devices operate in some form of network, with the goal of either increasing computational power, to perform distributed tasks or to simply exchange messages of some kind. The structure and implementations of such networks often depend on the sizes of devices. For example, for embedded devices such a network corresponds to connected, hard-wired micro-controllers. In contrast computers connect through local-area-networks over WiFi. With the advent of quantum computers approaching, researchers believe the next evolutionary step in the computing industry lies in the connection of quantum computing devices, with the ultimate goal to unleash new applications impossible by classical means. Such applications include quantum key distribution [1], secret sharing [2, 3], sensor networks [4, 5] but also distributed quantum computing [6, 7, 8].
@@
At the lowest level, quantum devices, in any quantum network [9, 10, 11, 12, 13], at any scale, are connected via physical channels. A Bell-state between nodes is equivalent to a channel due to the teleportation protocol [14], paving the way for sharing and manipulating entanglement [15] as an alternative to channels. In recent years, the vast majority of research on quantum networks has focused on the generation of Bell-states between two nodes in a network of Bell-states [16, 17, 18, 19, 20, 21, 22, 8, 13, 23, 24, 25]. This generation involves finding an optimal path [26] (in terms of some metrics) within the network and connecting the states along that path. Here, we introduce a flexible quantum data bus by establishing multiple Bell-states simultaneously in a network of any scale by utilizing pre-established 2D cluster states as resource, referred to as cluster states [27, 28] in the following, instead of a network of Bell-states as a resource for routing [29, 30].
@@
In our setting, devices may connect at any node in the cluster state, which is generated as a flexible resource state to fulfill different tasks on demand, even prior to network requests. The manipulation of resource states takes place solely by single-qubit measurements and single-qubit unitaries. At the heart of our proposal is the zipper-scheme, a measurement pattern among diagonal lines to establish a Bell-state within the cluster state. We remark that the zipper-scheme is based on the X-protocol introduced in [31]. Other works also study staircase-shaped measurement patterns on 2D cluster states, even demonstrating optimality with regards to the number of measurements [32, 33]. As we demonstrate, the zipper-scheme preserves the entanglement structure of the remaining cluster state when generating a Bell-state. Keeping the grid structure intact and usable turns out to be pivotal to extract further Bell-states from the cluster state. We utilize this remarkable property to establish multiple Bell-states between arbitrary nodes in a cluster state, including crossings and turns of the measurement paths. We note, however, that there exist some prohibited areas around the turning and end points, where Pauli Z measurements are necessary to cut out the Bell-state from its direct neighborhood. However, this is still in stark contrast to standard methods [34, 28, 35], where the connection between nodes is established by cutting a hole among the entire path. This method not only wastes resources, but also hinders the generation of further Bell-states. Our scheme enables to simultaneously create Bell-states along adjacent measurements paths, effectively constructing a quantum data bus in analogy to classical busses found for instance in mainboards. We compose a quantum data bus architecture in a modular fashion by combining elementary building blocks such as crossings, turns and fade-in/-outs. Our results have a multitude of applications from which we discuss its applicability ranging from large scale networks such as entanglement-based quantum networks  [29, 30, 36, 37] to local networks relying on a central unit [38, 39]. Furthermore, we believe that also small and integrated networks of quantum devices, similar to devices running in embedded environments like cars, benefit from a dynamic way of creating Bell-states on demand, as this removes the necessity to multiplex multiple connections over a single transmission channel like in classical, hard-wired controller networks. Diagonal routing also naturally extend to the multiparty case, like for example establishing GHZ states between multiple nodes within the cluster.
Setting.—
@@
We consider quantum networks at different scales with the goal of establishing entanglement between multiple pairs of nodes in parallel. Rather than generating Bell pairs directly, we consider an entanglement-based scenario where nodes share a pre-prepared multipartite entangled resource state which they manipulate by means of measurements to establish desired configurations in a flexible way on demand. Such an approach has the advantage that it reduces latency, and allows to optimize the entanglement topology [29, 40, 41]. On the contrary it requires a long time memory for storage. Here we consider a 2D cluster state as a resource, where each node holds a single qubit of the cluster state. Such 2D cluster states serve as universal resource in measurement-based quantum computation (MBQC) [34, 28, 42], where information is processed by means of single qubit measurements only and arbitrary target states can be generated in this way. Here, in contrast, our sole goal is to establish multiple Bell pairs, which implies that considerations such as information flow [43, 44] do not apply, and additional methods and techniques are available. In particular, we are interested in establishing (multiple) Bell pairs between arbitrary nodes, independent of their location within the cluster state.
Preliminaries.—
Graph states [45, 46] are pure, multipartite entangled quantum states described by a classical graph $G=(V,E)$, where the vertex set $V$ and edge set $E$ correspond to qubits and binary Ising-type interaction between those, respectively. Formally, a graph state |$\ket{G}$ is uniquely defined as the eigenvector with $+1$ eigenvalues of the set of stabilizers
@@
for each vertex $a\in V$ with its corresponding neighbors $N_{a}\subseteq V$. The dynamic picture is an alternative mean to define graph states, where all qubits are initially prepared in the $\ket{+}$ state and according to the edge set controlled-phase gates are applied between two vertices.
@@
In this work we frequently use local complementation (LC), a specific local Clifford unitary, to transform a graph state. This unitary inverts the sub-graph induced by the neighbors of a qubit. Local Pauli Z measurements translate to vertex and edge erasure of the measured qubit. A Pauli Y measurement corresponds to a LC followed by a Pauli Z measurement on the qubit to measure. Local Pauli X measurements are a combination of LC on a specific neighbor qubit, a Pauli Y measurement on the qubit itself and again LC on the specific neighbor, see Ref. [46].
Zipper-scheme.—
@@
The zipper-scheme tackles the problem of connecting two nodes in a 2D cluster state via two diagonal lines, thereby generating a Bell-state. The scheme accomplishes this by performing Pauli X measurements along a staircase-shaped measurement path (as introduced in [31] as X-protocol), where two such paths are combined to connect any two points on the grid. For a direct diagonal connection, one of the Bell-state qubits is used as neighbor for LC, and Pauli Z measurements in the neighbourhood of the endpoints isolate the final Bell-state. The inset in Fig. 1 illustrates the result of the zipper-scheme after the orange qubits have been measured in the X basis along a diagonal, whereas the purple qubit is one of the Bell-state constituents and has been used as reference qubit for LC. We observe that the zipper-scheme restores the remaining cluster state structure, see the red edges in Fig. 1, up to holes at the endpoints. Note that some Pauli Z basis measurements are necessary to fully isolate the Bell-state though, see the yellow qubits in Fig. 1. We investigate the properties of the protocol in depth in the Appendix.
Pauli X measurement
Pauli Z measurement
Qubit measured in Pauli X basis
@@
Figure 1: Zipper-scheme (inset): The orange qubits are measured in the Pauli X basis such that the purple qubit is routed along the diagonal measurement path while the obtained edges between the red qubits restore the underlying cluster state. The qubits in the neighborhood of the purple qubit need to be removed by Pauli Z measurements leading to holes. Simultaneous Bell-state routing (main figure): We extract three Bell-states, purple, blue and turquoise, by applying the zipper-scheme along diagonals of orange qubits. Merely on turning and end points holes appear due to the yellow Pauli Z measurements to disconnect the Bell-state qubits from the remaining cluster state. The zipper-scheme enables crossing of paths except in the vicinity of end and turning points, and note that the paths are adjusted such that the measurement sequence is taken into account (first purple, then turquoise and blue).
@@
The zipper-scheme allows us to generate a Bell-state on a diagonal line within the cluster state. However, we observe, as explained in more details in the Appendix (note that routing a single Bell-state is a special case of routing parallel Bell-states), that it is also possible to use the very same scheme to route a Bell-state around corners (’L’-turn), upside down (’V’-turn), diagonal-to-straight or over a crossing with another Bell-state. These ingredients enable us to route Bell-states of arbitrary topology on the cluster state. In Fig. 1 we demonstrate an example in which three Bell-states among arbitrary nodes are generated. We emphasize that two points on a 2D lattice can be always connected by two (or more) diagonals as shown in the Fig. 1. Importantly, the zipper-scheme allows to establish crossings of Bell-states on the cluster state provided the crossings are sufficiently far away from the end or turning points. Note that we do not sketch the additional edges after applying the zipper-scheme to simplify readability, and furthermore, we have adjusted the measurement paths assuming that the purple path is measured first followed by the turquoise and blue one. The ability to support crossings has two advantages. First, it prevents potentially long measurement paths around obstacles in a cluster state, and second, it allows to have simultaneous request also in cases that require crossings due to resource constraints, impossible to achieve without them. Using a 2D cluster state of size $n\times n$ allows one to obtain $\mathcal{O}(n)$ parallel Bell pairs. This can be seen by observing that any two points in a 2D cluster are connected by diagonal paths of length at most $2n$, which also corresponds (up to few additional measurements around end- and turning points) to the number of qubits that need to be measured to establish a Bell pair. Since the underlying entanglement structure is still intact, in total $\mathcal{O}(n)$ Bell pairs can be generated generically. Notice that some configurations may not be possible, as small holes are cut in the 2D structure around end- and turning points.
Quantum data bus.—
@@
The next step is to extend the results of the zipper-scheme and building blocks to route multiple Bell-states along parallel lines and to which we refer to as a quantum data bus. We note that a quantum data bus is conceptually closely related to data busses in conventional computers used for connecting individual microchips with each other. A quantum data bus uses the entanglement of a cluster state in an efficient way to achieve that goal. In Fig. 2 we illustrate the main idea.
@@
Figure 2: Quantum data bus and its building blocks: The red lines represent parallel transport along diagonal lines, including crossings orthogonal to the diagonals. The black lines show a parallel transport, both vertical and horizontal. The blue and pink lines depict parallel ’L’- and ’V’-turns, whereas the green lines demonstrate a fade-in/-out of data lines. Finally, the yellow qubits depict qubits on which we need to perform Pauli Z measurements.
Similar when routing a single Bell-state we identify multiple building blocks for routing Bell-states in parallel. Five of the parallel building blocks naturally extend the building blocks for a single Bell-state to multiple parallel Bell-states (diagonal, crossing, ’L’-turn, ’V’-turn, diagonal-to-straight). For instance, from a cluster state of size $n\times n$, it is possible to generate $\mathcal{O}(n)$ parallel Bell-states as part of a crossing. The required number of qubits to turn $n$ data lines in an ’L’-turn around the corner is $n\times n$. For ’V’-turns and the building block that turns diagonal Bell-states into a parallel Bell-states, the number of data lines $n$ of the building block determines the minimum, necessary length of the zipper-scheme, corresponding to $2n$ for the first data line. Therefore, it requires in total $n\times n$ measurements.
@@
A new building block corresponds to parallel transport among a straight line. We achieve this kind of transport by applying the zipper-scheme repeatedly. Specifically, first one applies the zipper-scheme from the first Bell-state line towards the last line. This effectively routes the first Bell-state diagonally to the last line. Next, one applies the same scheme to the second Bell-state, moving it to the second last line. Repeatedly applying this pattern results in a total permutation of all lines. Performing the same scheme another time inverts the permutation, resulting in straight lines again. However, it should be noted that additional Pauli Z measurements are necessary to isolate the established data bus. A single qubit line is kept between the two permutation measurements, because the zipper-scheme requires to have a cluster state. In total it requires a block of length $2n$ qubits to transport $n$ data lines in parallel, which induces a fixed, minimal block size. Notice, though that there is no distance between the data lines within the data bus, in stark contrast to standard data transport schemes in MBQC [34, 28, 35] which require to isolate a path, resulting in a distance of one between two data lines, i.e. half the capacity as in our scheme.
@@
Another new building block allows to fade-in or -out data to the data bus. It uses Pauli Z and X measurements to separate and merge the data lines, respectively. For adding a single data line we require one Pauli Z and one Pauli X measurement to cut and merge the line into the data bus.
@@
We discuss all building blocks and its measurements in the Appendix in more detail. We also point out that diagonal routing straightforwardly extends to GHZ states, as described in the Appendix. Essentially, one can connect several qubits to some central one via different (diagonal) paths, thereby establishing a GHZ state.
Applications.—

### v3

# Flexible quantum data bus for quantum networks

@@ Abstract
We consider multipath generation of Bell states in quantum networks, where a preprepared multipartite entangled two-dimensional cluster state serves as a resource to perform different tasks on demand. We show how to achieve parallel connections between multiple, freely chosen groups of parties by performing appropriate local measurements along a diagonal, staircase-shaped path on a two-dimensional cluster state. Remarkably, our measurement scheme preserves the entanglement structure of the cluster state such that the remaining state is again a two-dimensional cluster state. We demonstrate strategies for generating crossing, turning, and merging of multiple measurement lines along the two-dimensional cluster state. The results apply to local area as well as to long-distance networks.
@@ I Introduction
Connecting devices constitutes the foundation of modern society. All devices operate in some form of network, with the goal of either increasing computational power, to perform distributed tasks or to simply exchange messages of some kind. The structure and implementations of such networks often depend on the sizes of devices. For example, for embedded devices such a network corresponds to connected, hard-wired microcontrollers. In contrast, computers connect through local-area-networks over Wi-Fi. With the advent of quantum computers approaching, researchers believe that connecting quantum devices via quantum networks is a crucial step to unlock new applications which do not have a classical counterpart [1, 2]. Such applications include quantum key distribution [3], secret sharing [4, 5], sensor networks [6, 7] but also distributed quantum computing [8, 9, 10].
@@
At the lowest level, quantum devices, in any quantum network [11, 12, 13, 14, 15], at any scale, are connected via physical channels. A Bell state between nodes is equivalent to a channel due to the teleportation protocol [16], paving the way for sharing and manipulating entanglement [17] as an alternative to channels. In recent years, the vast majority of research on quantum networks has focused on the generation of Bell states between two nodes in a network of Bell states [18, 19, 20, 21], the routing of quantum information carriers [22, 23, 24, 25, 26, 27] in quantum networks. The generation of Bell states involves finding an optimal path [23] (in terms of some metrics) within the network and connecting the states along that path. Here, we introduce a flexible quantum data bus by establishing multiple Bell states simultaneously in a network of any scale by utilizing pre-established two-dimensional (2D) cluster states as a resource, referred to as cluster states [28, 29] in the following, instead of a network of Bell states as a resource for routing [30, 31].
In our setting, we assume that each network node stores only a single qubit. The nodes share a two-dimensional cluster state, which has a grid structure, and interconnects the nodes and devices, respectively. This grid-state serves as a flexible network state to fulfill different tasks on demand, and can be established even prior to network requests. The manipulation of network states takes place solely by single-qubit measurements and single-qubit unitaries. This falls within the general spirit of entanglement-based quantum networks with preshared entanglement [30, 31]. Such an approach offers not only high flexibility [32], but also avoids latencies.
In this work we report the following two main results:
  1. (i)
    A zipper scheme to establish a connection between two nodes within a two-dimensional cluster state, which restores the structure of the remaining two-dimensional cluster state.
  2. (ii)
    A quantum data bus that allows to generate multiple Bell states simultaneously by measuring along parallel measurement lines along the two-dimensional cluster state.
@@
At the heart of our proposal is the zipper scheme (i), a measurement pattern among diagonal lines to establish a Bell state within the grid of the two-dimensional cluster state. We remark that the zipper scheme is based on the X protocol introduced in Ref. [33]. Other works also study staircase-shaped measurement patterns on 2D cluster states, even demonstrating optimality with regards to the number of measurements [34, 35]. As we demonstrate, the zipper scheme preserves the entanglement structure of the remaining cluster state when generating a Bell state. Keeping the grid structure intact and usable turns out to be pivotal to extract further Bell states from the cluster state. We utilize this remarkable property to establish multiple Bell states between arbitrary nodes in a cluster state, including crossings and turns of the measurement paths. We note, however, that there exist some prohibited areas around the turning and endpoints, where Pauli $Z$ measurements are necessary to cut out the Bell state from its direct neighborhood. However, this is still in stark contrast to standard methods [36, 29, 37], where the connection between nodes is established by cutting a hole among the entire path. This method not only wastes resources, but also hinders the generation of further Bell states. Our scheme enables to simultaneously create Bell states along adjacent measurements paths, effectively constructing a quantum data bus (ii) in analogy to classical buses found for instance in mainboards. We compose a quantum data bus architecture in a modular fashion by combining elementary building blocks such as crossings, turns and merging or splitting of multiqubit measurement lines in a cluster state. Our results have a multitude of applications from which we discuss its applicability ranging from large-scale networks such as entanglement-based quantum networks [30, 31, 38, 39] to local networks relying on a central unit [40, 41]. Furthermore, we believe that also small and integrated networks of quantum devices, similar to devices running in embedded environments like cars, benefit from a dynamic way of creating Bell states on demand, as this removes the necessity to multiplex multiple connections over a single transmission channel like in classical, hardwired controller networks. Diagonal routing also naturally extend to the multiparty case, like for example establishing Greenberger-Horne-Zeilinger (GHZ) states between multiple nodes within the cluster.
This work is organized as follows. In section II we introduce our network setting based on entangled network states. In particular, we utilize graph states and their local manipulations to obtain the network requests, as we explain in section III. Subsequently, we present our main finding, the zipper scheme in section IV, and show how we can use these result to build a modular quantum data bus in section V. In section VI we focus on three applications of quantum networks which profit from our findings. Finally, we conclude our work and point out future research directions in section VII.
II Setting
@@
We consider quantum networks at different scales with the goal of establishing entanglement between multiple pairs of nodes in parallel. Rather than generating Bell pairs directly, we consider an entanglement-based network [30, 31] where nodes share a preprepared multipartite entangled network state, which is generated during idle times of the network. In this work we suggest a flat hierarchy among the nodes. In particular, we assume that each node holds one qubit of the cluster state, and thus all nodes have the same functionality and privileges. The nodes manipulate the network state solely by local unitaries and single-qubit measurements to establish the desired target state in a flexible way on demand. This approach does not require to send any quantum information carriers when the network request arrives and consequently reduces the latency to serve requests. Moreover, the network state can be optimized with respect to the connectivity demands of the network [32, 42] such as adjusting to local connectivity demands of client groups. On the contrary entanglement-based networks require a long time memory for storage. Here we consider a 2D cluster state as a resource, where each node holds a single qubit of the cluster state. Such 2D cluster states serve as universal resource in measurement-based quantum computation (MBQC) [36, 29, 43], where information is processed by means of single-qubit measurements only and arbitrary target states can be generated in this way. Here, in contrast, our sole goal is to establish multiple Bell pairs, which implies that considerations such as information flow [44, 45] do not apply, and additional methods and techniques are available. In particular, we are interested in establishing (multiple) Bell pairs between arbitrary nodes, independent of their location within the cluster state.
III Preliminaries
[…]
@@ III Preliminaries
[…]
@@
In this work we frequently use local complementation (LC), a specific local Clifford unitary, to transform a graph state. This unitary inverts the subgraph induced by the neighbors of a qubit. Local Pauli $Z$ measurements translate to vertex and edge erasure of the measured qubit. A Pauli $Y$ measurement corresponds to a LC followed by a Pauli $Z$ measurement on the qubit to measure. Local Pauli $X$ measurements are a combination of LC on a specific neighbor qubit, a Pauli $Y$ measurement on the qubit itself and again LC on the specific neighbor, see Ref. [47].
IV Zipper scheme
@@
The zipper scheme tackles the problem of connecting two nodes in a 2D cluster state via two diagonal lines, thereby generating a Bell state. The scheme accomplishes this by performing Pauli $X$ measurements along a staircase-shaped measurement path (as introduced in [33] as $X$ protocol), where two such paths are combined to connect any two points on the grid. For a direct diagonal connection, one of the Bell state qubits is used as neighbor for LC, and Pauli $Z$ measurements in the neighborhood of the endpoints isolate the final Bell state. The inset in Fig. 1 illustrates the result of the zipper scheme after the orange qubits have been measured in the $X$ basis along a diagonal, whereas the purple qubit is one of the Bell state constituents and has been used as reference qubit for LC. We observe that measuring the orange qubits along the diagonal, purple path introduces the red edges between the red qubits on each side of the measurement path, these edges restore the underlying 2D cluster state structure. However, some holes are caused by Pauli $Z$ measurements in the direct neighborhood of end- and turning points, which are necessary to fully isolate the Bell state, the yellow qubits in Fig. 1 correspond to this measurements.
@@
Figure 1: Zipper scheme (inset): The orange qubits are measured in the Pauli $X$ basis such that the purple qubit is routed along the diagonal measurement path while the obtained edges between the red qubits restore the underlying cluster state. The qubits in the neighborhood of the purple qubit need to be removed by Pauli $Z$ measurements leading to holes. Simultaneous Bell state routing (main figure): We extract three Bell states, purple, blue and turquoise, by applying the zipper scheme along diagonals of orange qubits. Merely on turning and endpoints holes appear due to the yellow Pauli $Z$ measurements to disconnect the Bell state qubits from the remaining cluster state. The zipper scheme enables crossing of paths except in the vicinity of end and turning points, and note that the paths are adjusted such that the measurement sequence is taken into account (first purple, then turquoise and blue).
@@
The zipper scheme allows us to generate a Bell state on a diagonal line within the cluster state. However, we observe, as explained in more details in the Appendix (note that routing a single Bell state is a special case of routing parallel Bell states), that it is also possible to use the very same scheme to route a Bell state around corners ('L'-turn), upside down ('V'-turn), diagonal-to-straight or over a crossing with another Bell state. These ingredients enable us to route measurement lines for Bell states of arbitrary topology on the cluster state. In Fig. 1 we demonstrate an example in which three Bell states among arbitrary nodes are generated. We emphasize that two points on a 2D lattice can be always connected by two (or more) diagonals as shown in the Fig. 1. Importantly, the zipper scheme allows us to establish crossings of Bell states on the cluster state provided that the crossings do not involve on qubits that are direct neighbors (yellow qubits to be measured in $Z$ basis) of the end or turning points.
Note that we do not sketch the additional edges after applying the zipper scheme to simplify readability, and furthermore, we have adjusted the measurement paths assuming that the purple path is measured first followed by the turquoise and blue one. The ability to support crossings has two advantages. First, it prevents potentially long measurement paths around obstacles in a cluster state, and second, it allows to have simultaneous request also in cases that require crossings due to resource constraints, impossible to achieve without them. Using a 2D cluster state of size $n\times n$ allows one to obtain $\mathcal{O}(n)$ parallel Bell pairs. This can be seen by observing that any two points in a 2D cluster are connected by diagonal paths of length at most $2n$, which also corresponds (up to few additional measurements around end- and turning points) to the number of qubits that need to be measured to establish a Bell pair. Since the underlying entanglement structure is still intact, in total $\mathcal{O}(n)$ Bell pairs can be generated generically. Notice that some configurations may not be possible, as small holes appear in the two-dimensional cluster state from $Z$ measurements of qubits that directly neighbor end- and turning point qubits on the measurement path.
V Quantum data bus
@@
The next step is to extend the results of the zipper scheme and building blocks to generate multiple Bell states along parallel lines, which we refer to as a quantum data bus. We note that a quantum data bus is conceptually closely related to data buses in conventional computers used for connecting individual microchips with each other. First, our scheme generates multiple, spatially separated Bell pairs along parallel measurement lines, which is analogously found in classical data buses. The second main feature of a classical data bus is that it is bidirectional, which means that information can be sent in both directions, and our measurement scheme allows for that. A quantum data bus uses the entanglement of a cluster state in an efficient way to achieve that goal. In Fig. 2 we illustrate the main idea.
@@
Figure 2: Quantum data bus and its building blocks: The red lines represent parallel transport along diagonal lines, including crossings orthogonal to the diagonals. The black lines show a parallel transport, both vertical and horizontal. The blue and pink lines depict parallel 'L'- and 'V'-turns, whereas the green lines demonstrate a merging or splitting of data lines. Finally, the yellow qubits depict qubits on which we need to perform Pauli $Z$ measurements.
Similar when generating a single Bell state along a measurement line, we identify multiple building blocks for generating Bell states in parallel. We point out that the building blocks we discussed in the previous section extend naturally to parallel building blocks. Foremost, is the diagonal crossing of Bell states, which follows trivially from the state restoring properties of the zipper scheme, and we represent it by the red structure in Fig. 2. For instance, from a cluster state of size $n\times n$, it is possible to generate $\mathcal{O}(n)$ parallel Bell states as part of a crossing. The entanglement structure remains when one applies the zipper scheme, and this key feature gives rise for further building blocks such as the 'L'(blue blocks) and 'V'-turn (purple blocks) as well as the straight line measurement lines (black blocks), which we discuss partly here and in the Appendix E.
Here we begin with the 'L'-turn building block for the quantum data bus, which can be used to change the direction of the measurement lines from vertical to horizontal (and vice versa), which illustrate step-by-step in Fig. 3. We apply the zipper scheme (in inward direction) to the outermost data line until we diagonally reach the inner most data line, resulting in a perpendicular turn as illustrated by the leftmost blue line of Fig. 3. The cluster state preservation property of this first step enables us to apply the zipper scheme to turn the second most left qubit as shown in the third step of Fig. 3. Repeatedly applying this pattern results in a full turn of all data lines, as illustrated in the last step of Fig. 3. Note that the required number of qubits to turn $n$ data lines in an 'L'-turn around the corner is $n\times n$.
Figure 3: The figure shows a step-by-step guide for the measurement pattern to perform a perpendicular turn from vertical to horizontal in a 2D cluster state with four parallel data lines. The zipper scheme is applied from the left- to right-most qubits, where each application of the zipper scheme enables the next measurement. Note that the red arrows are only visual elements to better see the turn of data lines, and the arrows can be also reversed.
@@
A new building block corresponds to parallel transport among a straight line. We achieve this kind of transport by applying the zipper scheme repeatedly. We point out that one can understand this building block simply from the previous one, if we extend the zipper scheme in Fig. 3 to the yellow left most qubits by $X$ measuring the top qubits (indicated by the arrows). Specifically, first one applies the zipper scheme from the first Bell state line towards the last line. This effectively routes the first Bell state diagonally to the last line. Next, one applies the same scheme to the second Bell state, moving it to the second last line. Repeatedly applying this pattern results in a total permutation of all lines. Performing the same scheme another time inverts the permutation, resulting in straight lines again. However, it should be noted that additional Pauli $Z$ measurements are necessary to isolate the established data bus as shown by yellow dots and in Fig. 2 and Fig. 3. A single vertical qubit line is necessary to separate the two permutation measurements, because the zipper scheme requires be applied on a complete a cluster state. In total it requires a block of length $2n$ qubits to transport $n$ data lines in parallel, which induces a fixed, minimal block size. Notice, although that there is no distance between the data lines within the data bus, in stark contrast to standard data transport schemes in MBQC [36, 29, 37] which require to isolate a path, resulting in a distance of one between two data lines, i.e. half the capacity as in our scheme.
@@
Another new building block allows us to merge or split data lines to or from the data bus. In Fig. 2 the green structure represents a splitting of data lines, and it uses Pauli $Z$ (yellow dots) and $X$ measurements (green dots) to separate and merge the data lines, respectively. For adding a single data line we require one Pauli $Z$ and one Pauli $X$ measurement to cut and merge the line into the data bus.
@@
Similar approaches as presented apply for other building blocks as well, we refer to the Appendixes CE, and F for details. For 'V'-turns and the building block that turns diagonal Bell states into a parallel Bell states, the number of data lines $n$ of the building block determines the minimum, necessary length of the zipper scheme on the first data line as $2n$. Therefore, it requires in total $n\times n$ measurements. We also point out that diagonal routing straightforwardly extends to GHZ states, as described in the Appendix G. Essentially, one can connect several qubits to some central one via different (diagonal) paths, thereby establishing a GHZ state.
The size of the individual building blocks allows one to estimate the size of the required cluster state depending on the desired information transport.
VI Applications